# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ 5 Experiments
Table 1 summarizes the performance of GALAX and several competitive baselines across the full test dataset, with LUAD and BRCA representing the cancer types with the largest sample sizes in Target-QA. GALAX significantly outperforms all baselines on every evaluation metric. For example, it achieves an overall precision of 0.5472 and recall of 0.5332, exceeding the strongest baseline L3-FT (QA) + Omic in both metrics. The model also demonstrates strong performance across LUAD, BRCA, and other cancer-specific subsets, highlighting its robustness and generalizability. To further assess the quality of target prioritization, we evaluate Hit@10 and Hit@5, as shown in Table 2. GALAX achieves the highest accuracy across most evaluation settings, with an overall Hit@10 of 0.8815 and Hit@5 of 0.9249. Figure 4b-c illustrates the explainable subgraph generated for the lung cancer cell line ACH-000860. To further validate the biological relevance of the extracted subgraph, we performed functional enrichment analysis. The results reveal significant enrichment in cancer-associated signaling pathways, including the cancer pathway WP5434 and EGFR-related receptor signaling pathways such as WP138 and WP3680, as cataloged in WikiPathways (Agrawal et al., 2024). Notably, EGFR a well-established therapeutic target in NSCLC (Steuer & Ramalingam, 2015) appears in five enriched terms, together with PTK2 and WNT16 which are known to regulate invasion, epithelial mesenchymal transition, and therapeutic resistance (Tong et al., 2019; Sun et al., 2012). The observed pathway enrichment provides strong biological support for the relevance of the selected targets in lung cancer. Additional disease enrichment using the GAD DISEASE database further supports this conclusion (Sherman et al., 2022), with lung cancer identified as the top associated disease term with a p value of 0.0022, involving GSTM3, APAF1, NOD2, MLLT3, GC and EGFR. Full enrichment details across all cancer types are included in the Appendix  D.4.
@@ 6 Conclusion
We present GALAX, a graph-augmented language model that integrates large language models with pretrained graph neural networks under the guidance of a Graph Process Reward Model (GPRM) to enable target prioritization through explainable subgraph reasoning. GALAX constructs subgraphs in a step-wise manner via reinforcement learning, using pretrained GNNs to evaluate the biological plausibility of each step—removing the need for intermediate annotations. This framework supports structured, clinically grounded reasoning over multi-omic data. To facilitate evaluation, we introduce Target-QA, a benchmark dataset integrating multi-omics and CRISPR outcomes. GALAX achieves performance gains over baselines across most evaluation metrics on Target-QA.
@@
This dataset is derived from the Broad Institute’s Cancer Dependency Map (DepMap [^3] DepMap: https://depmap.org/portal/) and will be released strictly for non-commercial, internal research and academic use, consistent with DepMap’s Terms of Use. We do not redistribute original DepMap files; instead, we provide derived, non-identifiable annotations and processing scripts/pointers so users can obtain the source data directly from DepMap after accepting its terms. The dataset is not intended for clinical applications and must not be used for any Commercial Use (e.g., direct sale, incorporation into a product, or training/developing/enhancing ML/AI models beyond internal academic research). Users agree to acknowledge DepMap and the Broad Institute using the acknowledgement wording specified by DepMap, and to respect any third-party rights that may attach to the underlying data. Users must preserve confidentiality and refrain from any re-identification attempts. This statement summarizes our compliance posture and does not constitute legal advice; users are responsible for ensuring their own compliance with DepMap’s Terms and applicable policies.
@@
We release the source code of GALAX [^4] GitHub: https://github.com/FuhaiLiAiLab/GALAX in GitHub, together with preprocessing pipelines, thereby enabling reproduction of Target-QA datasets and reported experiments. Target-QA [^5] Hugging Face: https://huggingface.co/datasets/FuhaiLiAiLab/Target-QA dataset is also available in Huggingface.
@@ D.1 Prompt and Context Design for Finetuning
Same as aforementioned prompt design shown in Table  6 , we construct the final-stage prompt $P_{n}^{\text{(final)}}$, which incorporates both the initial answering and the subgraph-based regulatory context, as detailed in Table 7. We perform five additional finetuning trials using this refined prompt format. The corresponding training are illustrated in Figure 10, where the best run (run 5) demonstrates smooth convergence and is selected as the final model checkpoint $f_{\text{final}}$ for generating the ultimate target predictions.
@@ D.2 Hyperparameters
To generate high-quality and biologically reasonable subgraph explanations, we implement a reinforcement-guided generator, denoted as $\pi(\cdot)$, which operates over sample-specific graph states and candidate sets. To enhance robustness under noisy or unstable generation dynamics, we introduce a retry-based mechanism that adaptively tunes key hyperparameters with multiple runs ($\Omega=6$, and $\omega$ denotes the current number of runs), dynamically adjusting the configuration to encourage exploration and improve convergence. Specifically, the number of training epochs is reduced from 5 to a minimum of 2, promoting faster reinitialization. Simultaneously, the learning rate is increased linearly from an initial value of $1\times 10^{-3}$ to $0.001\cdot(1+\omega)$ to escape local minima. The maximum number of nodes per graph is reduced from 200 to 100 in steps of 25. The maximum number of graph construction steps $i$ per rollout is similarly reduced from 50 to 20 in steps of 5. The retry mechanism enables efficient navigation of the search space while preserving biological plausibility through domain-specific priors embedded in the reward formulation.

### v2

@@ 5 Experiments
Tables 12 summarize the performance of GALAX and several competitive baselines on the full test dataset. GALAX outperforms all baselines on every metric, reaching an overall precision of 0.5472 and recall of 0.5332. To further assess target prioritization quality, we report Hit@10 and Hit@5 in Table 2, where GALAX again achieves the highest accuracy with an overall Hit@10 of 0.8815 and Hit@5 of 0.9249. We also replaced the backbone with Qwen2.5–7B–Instruct and observed similar performances on all metrics, indicating that GALAX maintains stable performance under backbone changes (See Appendix D.4 for full experiment results). We further tested generalization by forming three holdout sets in which the selected TCGA cancer types were excluded from training and treated as unseen during evaluation. Across all holdout sets, GALAX showed only modest performance declines, indicating that it preserves strong accuracy and generalizes reliably to previously unseen cancer types (Details are provided in Appendix  D.5). Beyond Target-QA, we evaluate GALAX on the pediatric cancer dataset from PedDep (dharia2021first), which offers multi-omic profiles and CRISPR-based targets for 31 tumor cell lines, using zero-shot inference due to its small sample size. Even under this setting, GALAX surpasses all baselines on all metrics, showing strong transfer to external dataset (Details are provided in Appendix D.6). Figure 4b-c illustrates the explainable subgraph generated for the lung cancer cell line ACH-000860. To further validate the biological relevance of the extracted subgraph, we performed functional enrichment analysis. The results reveal significant enrichment in cancer-associated signaling pathways, including the cancer pathway WP5434 and EGFR-related receptor signaling pathways such as WP138 and WP3680, as cataloged in WikiPathways (agrawal2024wikipathways). Notably, EGFR a well-established therapeutic target in NSCLC (steuer2015targeting) appears in five enriched terms, together with PTK2 and WNT16 which are known to regulate invasion, epithelial mesenchymal transition, and therapeutic resistance (tong2019protein; sun2012treatment). The observed pathway enrichment provides strong biological support for the relevance of the selected targets in lung cancer. Additional disease enrichment using the GAD DISEASE database further supports this conclusion (sherman2022david), with lung cancer identified as the top associated disease term with a p value of 0.0022, involving GSTM3, APAF1, NOD2, MLLT3, GC and EGFR. Full enrichment details across all cancer types are included in the Appendix E.1. In addition, human and LLM evaluation results in Appendix E.2 show that most generated subgraphs are biologically plausible.
Table 4: Overall performances for ablation studies on omic inputs and KG structure
Config Setting Recall $\uparrow$ Precision $\uparrow$ Hit@5 $\uparrow$ Hit@10 $\uparrow$
GALAX
E20 Drop $20\%$ edges $0.5061{\scriptstyle\pm 0.0268}$ $0.5362{\scriptstyle\pm 0.0059}$ $0.9005{\scriptstyle\pm 0.0128}$ $0.8741{\scriptstyle\pm 0.0060}$
E40 Drop $40\%$ edges $0.2871{\scriptstyle\pm 0.0060}$ $0.5079{\scriptstyle\pm 0.0039}$ $0.8762{\scriptstyle\pm 0.0138}$ $0.8471{\scriptstyle\pm 0.0090}$
E60 Drop $60\%$ edges $0.2753{\scriptstyle\pm 0.0020}$ $0.4961{\scriptstyle\pm 0.0059}$ $0.8635{\scriptstyle\pm 0.0084}$ $0.8307{\scriptstyle\pm 0.0033}$
E80 Drop $80\%$ edges $0.2775{\scriptstyle\pm 0.0134}$ $0.4943{\scriptstyle\pm 0.0177}$ $0.8434{\scriptstyle\pm 0.0524}$ $0.8185{\scriptstyle\pm 0.0494}$
N20 Drop $20\%$ nodes $0.2697{\scriptstyle\pm 0.0013}$ $0.5034{\scriptstyle\pm 0.0014}$ $0.8786{\scriptstyle\pm 0.0070}$ $0.8560{\scriptstyle\pm 0.0092}$
N40 Drop $40\%$ nodes $0.2675{\scriptstyle\pm 0.0033}$ $0.4901{\scriptstyle\pm 0.0035}$ $0.8878{\scriptstyle\pm 0.0066}$ $0.8503{\scriptstyle\pm 0.0048}$
N60 Drop $60\%$ nodes $0.2617{\scriptstyle\pm 0.0056}$ $0.4929{\scriptstyle\pm 0.0045}$ $0.8698{\scriptstyle\pm 0.0290}$ $0.8385{\scriptstyle\pm 0.0247}$
N80 Drop $80\%$ nodes $0.2653{\scriptstyle\pm 0.0032}$ $0.4825{\scriptstyle\pm 0.0090}$ $0.8341{\scriptstyle\pm 0.0105}$ $0.8103{\scriptstyle\pm 0.0134}$
Omic-M Remove epigenomic data $0.4810{\scriptstyle\pm 0.0137}$ $0.5163{\scriptstyle\pm 0.0086}$ $0.8857{\scriptstyle\pm 0.0145}$ $0.8614{\scriptstyle\pm 0.0115}$
Omic-G Remove genomic data $0.3121{\scriptstyle\pm 0.0052}$ $0.4277{\scriptstyle\pm 0.0056}$ $0.8550{\scriptstyle\pm 0.0037}$ $0.8402{\scriptstyle\pm 0.0033}$
Omic-T Remove transcriptomic data $0.3377{\scriptstyle\pm 0.0016}$ $0.4065{\scriptstyle\pm 0.0042}$ $0.8720{\scriptstyle\pm 0.0037}$ $0.8672{\scriptstyle\pm 0.0037}$
Omic-P Remove proteomic data $0.3347{\scriptstyle\pm 0.0013}$ $0.3980{\scriptstyle\pm 0.0058}$ $0.8540{\scriptstyle\pm 0.0138}$ $0.8466{\scriptstyle\pm 0.0040}$
Omic-All Remove all omics $0.3024{\scriptstyle\pm 0.0032}$ $0.3793{\scriptstyle\pm 0.0019}$ $0.8237{\scriptstyle\pm 0.0066}$ $0.7967{\scriptstyle\pm 0.0040}$
GALAX Original $0.5332{\scriptstyle\pm 0.0031}$ $0.5472{\scriptstyle\pm 0.0053}$ $0.9249{\scriptstyle\pm 0.0048}$ $0.8815{\scriptstyle\pm 0.0033}$
L3-FT(QA) + Omic + KG
E20 Drop $20\%$ edges $0.4599{\scriptstyle\pm 0.0820}$ $0.5214{\scriptstyle\pm 0.0036}$ $0.8587{\scriptstyle\pm 0.0157}$ $0.8373{\scriptstyle\pm 0.0258}$
E40 Drop $40\%$ edges $0.2742{\scriptstyle\pm 0.0090}$ $0.5064{\scriptstyle\pm 0.0099}$ $0.8429{\scriptstyle\pm 0.0173}$ $0.8226{\scriptstyle\pm 0.0169}$
E60 Drop $60\%$ edges $0.2676{\scriptstyle\pm 0.0062}$ $0.4991{\scriptstyle\pm 0.0099}$ $0.8296{\scriptstyle\pm 0.0305}$ $0.8111{\scriptstyle\pm 0.0175}$
E80 Drop $80\%$ edges $0.2611{\scriptstyle\pm 0.0074}$ $0.4880{\scriptstyle\pm 0.0086}$ $0.8254{\scriptstyle\pm 0.0361}$ $0.8063{\scriptstyle\pm 0.0370}$
N20 Drop $20\%$ nodes $0.2662{\scriptstyle\pm 0.0059}$ $0.4916{\scriptstyle\pm 0.0024}$ $0.8434{\scriptstyle\pm 0.0186}$ $0.8222{\scriptstyle\pm 0.0193}$
N40 Drop $40\%$ nodes $0.2658{\scriptstyle\pm 0.0049}$ $0.4838{\scriptstyle\pm 0.0169}$ $0.8709{\scriptstyle\pm 0.0422}$ $0.8339{\scriptstyle\pm 0.0335}$
N60 Drop $60\%$ nodes $0.2648{\scriptstyle\pm 0.0000}$ $0.4742{\scriptstyle\pm 0.0150}$ $0.7857{\scriptstyle\pm 0.0247}$ $0.7690{\scriptstyle\pm 0.0303}$
N80 Drop $80\%$ nodes $0.2689{\scriptstyle\pm 0.0042}$ $0.4722{\scriptstyle\pm 0.0063}$ $0.7111{\scriptstyle\pm 0.0055}$ $0.6974{\scriptstyle\pm 0.0111}$
Omic-M Remove epigenomic data $0.4602{\scriptstyle\pm 0.0361}$ $0.4878{\scriptstyle\pm 0.0256}$ $0.8794{\scriptstyle\pm 0.0055}$ $0.8466{\scriptstyle\pm 0.0142}$
Omic-G Remove genomic data $0.3213{\scriptstyle\pm 0.0047}$ $0.3962{\scriptstyle\pm 0.0093}$ $0.8455{\scriptstyle\pm 0.0073}$ $0.8349{\scriptstyle\pm 0.0136}$
Omic-T Remove transcriptomic data $0.3244{\scriptstyle\pm 0.0034}$ $0.3996{\scriptstyle\pm 0.0087}$ $0.8550{\scriptstyle\pm 0.0073}$ $0.8381{\scriptstyle\pm 0.0097}$
Omic-P Remove proteomic data $0.3266{\scriptstyle\pm 0.0012}$ $0.3872{\scriptstyle\pm 0.0032}$ $0.8497{\scriptstyle\pm 0.0048}$ $0.8265{\scriptstyle\pm 0.0056}$
Omic-All Remove all omics $0.2669{\scriptstyle\pm 0.0075}$ $0.3577{\scriptstyle\pm 0.0093}$ $0.7830{\scriptstyle\pm 0.0182}$ $0.7538{\scriptstyle\pm 0.0174}$
L3-FT(QA)+Omic+KG Original $0.4908{\scriptstyle\pm 0.0402}$ $0.5185{\scriptstyle\pm 0.0240}$ $0.8794{\scriptstyle\pm 0.0114}$ $0.8529{\scriptstyle\pm 0.0153}$
Ablation Studies
As shown in Table 4, we evaluated GALAX under systematic perturbations to omic inputs and KG structure. The KG provides both the guidance for graph generation and the retrieved subgraphs that supply biological context to the language model, so removing portions of the KG naturally reduces overall performance. Under KG deletions, GALAX remained stronger than the baseline across all edge-removal levels because node attributes and pretrained GNN embeddings enabled the model to assemble meaningful subnetworks from sparse structure. Node deletion was more destructive since removing 20% of nodes removed about 35% of edges, but GALAX maintained more stable Hit@5 and Hit@10 scores than the baseline across all deletion levels. Removing epigenomic data produced the smallest decline due to methylation saturation, while removing genomic, transcriptomic, or proteomic signals caused larger drops, yet GALAX still outperformed the L3-FT(QA)+Omic+KG baseline even when all omics were removed. Overall, these findings show that GALAX generalizes well under shifts in omic distributions and reduced KG connectivity.
@@ 6 Conclusion
We present GALAX, a graph-augmented language model that unifies numerical multi-omic evidence, literature-scale textual information, and stepwise graph construction under reinforcement learning with biologically grounded supervision through a Graph Process Reward Model (GRPM), which scores intermediate subgraphs for biological plausibility and cancer relevance and guides the system to generate patient-specific mechanistic subgraph for target prioritization without explicit labels. To facilitate evaluation, we introduce Target-QA, a benchmark combining multi-omic data, CRISPR outcomes, and graph knowledge for target discovery. GALAX consistently outperforms baseline models on this dataset.
@@
This dataset is derived from the Broad Institute’s Cancer Dependency Map (DepMap [^3] https://depmap.org/portal/) and will be released strictly for non-commercial, internal research and academic use, consistent with DepMap’s Terms of Use. We do not redistribute original DepMap files; instead, we provide derived, non-identifiable annotations and processing scripts/pointers so users can obtain the source data directly from DepMap after accepting its terms. The dataset is not intended for clinical applications and must not be used for any Commercial Use (e.g., direct sale, incorporation into a product, or training/developing/enhancing ML/AI models beyond internal academic research). Users agree to acknowledge DepMap and the Broad Institute using the acknowledgement wording specified by DepMap, and to respect any third-party rights that may attach to the underlying data. Users must preserve confidentiality and refrain from any re-identification attempts. This statement summarizes our compliance posture and does not constitute legal advice; users are responsible for ensuring their own compliance with DepMap’s Terms and applicable policies.
@@
We release the source code in the supplementary materials, together with preprocessing pipelines, thereby enabling reproduction of Target-QA datasets and reported experiments.
Acknowledgments
This research was partially supported by NLM 1R01LM013902-01A1.
@@ B.2 DepMap Data Preprocessing
Table 5: TCGA cancer type codes and their full names.
TCGA Code Full Name TCGA Code Full Name
Overall Average Overall Average MB Medulloblastoma
LUAD Lung Adenocarcinoma ALL Acute Lymphoblastic Leukemia
BRCA Breast Invasive Carcinoma LGG Brain Lower Grade Glioma
COAD/READ Colon/Rectum Adenocarcinoma NB Neuroblastoma
PAAD Pancreatic Adenocarcinoma MESO Mesothelioma
GBM Glioblastoma Multiforme LIHC Liver Hepatocellular Carcinoma
SARC Sarcoma LAML Acute Myeloid Leukemia
OV Ovarian Serous Cystadenocarcinoma DLBC Lymphoid Neoplasm Diffuse Large B-cell Lymphoma
SKCM Skin Cutaneous Melanoma $M$ Multiple Myeloma
ESCA Esophageal Carcinoma KIRC Kidney Renal Clear Cell Carcinoma
SCLC Small Cell Lung Cancer THCA Thyroid Carcinoma
HNSC Head and Neck Squamous Cell Carcinoma BLCA Bladder Urothelial Carcinoma
LUSC Lung Squamous Cell Carcinoma UCEC Uterine Corpus Endometrial Carcinoma
STAD Stomach Adenocarcinoma PRAD Prostate Adenocarcinoma
@@ D.1 Prompt and Context Design for Finetuning
Same as aforementioned prompt design shown in Table  8 , we construct the final-stage prompt $P_{n}^{\text{(final)}}$, which incorporates both the initial answering and the subgraph-based regulatory context, as detailed in Table 9. We perform five additional finetuning trials using this refined prompt format. The corresponding training are illustrated in Figure 10, where the best run (run 5) demonstrates smooth convergence and is selected as the final model checkpoint $f_{\text{final}}$ for generating the ultimate target predictions.
@@ D.2 Hyperparameters
To generate high-quality and biologically reasonable subgraph explanations, we implement a reinforcement-guided generator, denoted as $\pi(\cdot)$, which operates over sample-specific graph states and candidate sets. To enhance robustness under noisy or unstable generation dynamics, we introduce a retry-based mechanism that adaptively tunes key hyperparameters with multiple runs ($\Psi=6$, and $\psi$ denotes the current number of runs), dynamically adjusting the configuration to encourage exploration and improve convergence. Specifically, the number of training epochs is reduced from 5 to a minimum of 2, promoting faster reinitialization. Simultaneously, the learning rate is increased linearly from an initial value of $1\times 10^{-3}$ to $0.001\cdot(1+\psi)$ to escape local minima. The maximum number of nodes per graph is reduced from 200 to 100 in steps of 25. The maximum number of graph construction steps $i$ per rollout is similarly reduced from 50 to 20 in steps of 5. The retry mechanism enables efficient navigation of the search space while preserving biological plausibility through domain-specific priors embedded in the reward formulation. To further study the effect of sparse versus dense rewards, we vary the reward calculation frequency by introducing a step interval parameter $\mathbf{s}$. Specifically, when $\mathbf{s}=2$, we compute the reward only every 2 steps during graph generation. Larger values of $\mathbf{s}$ produce sparser, more delayed reward signals, while smaller values (approaching $\mathbf{s}=1$) provide denser, more immediate feedback. As shown in Table 10, GALAX demonstrates consistent stability across all reward calculation intervals. Notably, performance improves as rewards become denser (smaller $\mathbf{s}$), with optimal results achieved at $\mathbf{s}=1$, the default GALAX configuration. This trend indicates that the graph generator benefits most from immediate biological feedback provided at each step by the graph process reward model. In contrast, less frequent reward calculations (larger $\mathbf{s}$) introduce noisier, longer-horizon estimates that dilute the signal quality. Importantly, even under highly sparse reward conditions (e.g., $\mathbf{s}=10$), GALAX consistently outperforms the baseline L3-FT(QA) + Omics + KG, demonstrating robustness to reward sparsity.
Table 10: Ablation study on reward density controlled by step interval $\mathbf{s}$
Reward Interval ($\mathbf{s}$) Precision $\uparrow$ Recall $\uparrow$ F1 Score $\uparrow$ Jaccard $\uparrow$ Hit@5 $\uparrow$ Hit@10 $\uparrow$
10 0.5312 $\pm$ 0.0044 0.5163 $\pm$ 0.0019 0.5232 $\pm$ 0.0031 0.3582 $\pm$ 0.0035 0.8996 $\pm$ 0.0050 0.8753 $\pm$ 0.0044
8 0.5347 $\pm$ 0.0062 0.5224 $\pm$ 0.0043 0.5284 $\pm$ 0.0054 0.3619 $\pm$ 0.0055 0.9063 $\pm$ 0.0079 0.8771 $\pm$ 0.0043
6 0.5398 $\pm$ 0.0051 0.5218 $\pm$ 0.0036 0.5316 $\pm$ 0.0044 0.3657 $\pm$ 0.0044 0.9024 $\pm$ 0.0079 0.8794 $\pm$ 0.0040
4 0.5389 $\pm$ 0.0054 0.5278 $\pm$ 0.0025 0.5302 $\pm$ 0.0041 0.3673 $\pm$ 0.0046 0.9137 $\pm$ 0.0056 0.8789 $\pm$ 0.0038
2 0.5463 $\pm$ 0.0057 0.5296 $\pm$ 0.0032 0.5384 $\pm$ 0.0040 0.3718 $\pm$ 0.0043 0.9218 $\pm$ 0.0048 0.8806 $\pm$ 0.0033
1 (GALAX) 0.5472 $\pm$ 0.0053 0.5332 $\pm$ 0.0031 0.5399 $\pm$ 0.0041 0.3726 $\pm$ 0.0037 0.9249 $\pm$ 0.0048 0.8815 $\pm$ 0.0033
L3-FT(QA) + Omics + KG 0.5185 $\pm$ 0.0240 0.4908 $\pm$ 0.0402 0.5038 $\pm$ 0.0327 0.3393 $\pm$ 0.0298 0.8794 $\pm$ 0.0114 0.8529 $\pm$ 0.0153
Furthermore, GALAX employs four RL components to collectively prevent reward hacking. First, the frozen graph oncogenicity classifier (Graph) derives reward signals from a pre-trained, frozen foundation model rather than a co-trained reward model, thereby eliminating co-evolution exploits. Second, the schema-based rule term (Rules) validates each action against biomedical knowledge graph ground truth to ensure that proposed edges conform to established biological constraints. Third, the rollout-based future reward (Rollout) evaluates each action based on its long-term consequences, preventing the generator from exploiting local reward irregularities or pursuing myopic gains. Fourth, stepwise quality gating (Gating) rejects actions yielding negative cumulative reward, ensuring that only biologically valid steps—those satisfying both schema rules and oncogenicity criteria—are retained. Together, these four components constitute a multi-faceted defense against reward hacking. To validate this design, we conduct an ablation study by systematically removing each component. As shown in Table 11, the robustness of GALAX does not stem from any single component but rather from the interplay of all four mechanisms. The rollout and gating modules help avoid short-sighted decisions and filter out invalid reasoning paths, while the graph supervisor and schema rules provide the underlying constraints that encourage the model to generate biologically plausible subgraphs. Notably, removing either the graph supervisor or the schema rules leads to consistent performance drops, suggesting that the frozen graph oncogenicity classifier offers reliable biological guidance, and that the schema constraints effectively block structurally incorrect actions. In sum, these mechanisms work collectively to push the generator toward producing biologically plausible subgraphs rather than outputs that score well but lack plausibility, making our RL considerably more robust than prior approaches.
Table 11: Ablation study of reward components
Model Variant Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ Jaccard $\uparrow$ Hit@5 $\uparrow$ Hit@10 $\uparrow$
L3-FT(QA) + Omics $0.5250\scriptstyle\pm 0.0282$ $0.4959\scriptstyle\pm 0.0435$ $0.5094\scriptstyle\pm 0.0365$ $0.3449\scriptstyle\pm 0.0338$ $0.8889\scriptstyle\pm 0.0168$ $0.8693\scriptstyle\pm 0.0157$
L3-FT(QA) + Omics + KG $0.5185\scriptstyle\pm 0.0240$ $0.4908\scriptstyle\pm 0.0402$ $0.5038\scriptstyle\pm 0.0327$ $0.3393\scriptstyle\pm 0.0298$ $0.8794\scriptstyle\pm 0.0114$ $0.8529\scriptstyle\pm 0.0153$
GALAX w/o Graph $0.5314\scriptstyle\pm 0.0045$ $0.5275\scriptstyle\pm 0.0077$ $0.5336\scriptstyle\pm 0.0053$ $0.3680\scriptstyle\pm 0.0043$ $0.9048\scriptstyle\pm 0.0084$ $0.8741\scriptstyle\pm 0.0056$
GALAX w/o Rules $0.5305\scriptstyle\pm 0.0074$ $0.5154\scriptstyle\pm 0.0115$ $0.5221\scriptstyle\pm 0.0098$ $0.3570\scriptstyle\pm 0.0078$ $0.8974\scriptstyle\pm 0.0073$ $0.8746\scriptstyle\pm 0.0042$
GALAX w/o Rollout $0.5362\scriptstyle\pm 0.0060$ $0.5196\scriptstyle\pm 0.0052$ $0.5271\scriptstyle\pm 0.0046$ $0.3628\scriptstyle\pm 0.0043$ $0.9067\scriptstyle\pm 0.0062$ $0.8769\scriptstyle\pm 0.0055$
GALAX w/o Gating $0.5387\scriptstyle\pm 0.0050$ $0.5243\scriptstyle\pm 0.0042$ $0.5308\scriptstyle\pm 0.0024$ $0.3651\scriptstyle\pm 0.0032$ $0.9118\scriptstyle\pm 0.0056$ $0.8782\scriptstyle\pm 0.0046$
GALAX (Full Model) $\mathbf{0.5472}\scriptstyle\pm\mathbf{0.0053}$ $\mathbf{0.5332}\scriptstyle\pm\mathbf{0.0031}$ $\mathbf{0.5399}\scriptstyle\pm\mathbf{0.0041}$ $\mathbf{0.3726}\scriptstyle\pm\mathbf{0.0037}$ $\mathbf{0.9249}\scriptstyle\pm\mathbf{0.0048}$ $\mathbf{0.8815}\scriptstyle\pm\mathbf{0.0033}$